OpenStudy (anonymous):

Integrate (2x +1)/(3x -1) I believe it can somehow be turned into a 1/x form, but my algebra is horrendous.

OpenStudy (phi):

try dividing 3x-1 into 2x+1 you get 2/3 + 5/(3(3x-1))

OpenStudy (anonymous):

\[\int\limits \frac{ 2 }{ 3 } + \frac{ 5 }{ 3(3x-1) }\] Is this what you're saying I then get?